1.) A pipe can fill a pool in 6 hours. A smaller pipe can fill the same pool in 8 hours. The empty pool needed to be filled as soon as possible, so the caretaker opened both pipes. However, he also mistakenly opened the drain. If the drain can empty the pool in 12 hours, how long will it take to fill the pool with both pipes and drain open?

Answer by stanbon(75887)

A pipe can fill a pool in 6 hours. A smaller pipe can fill the same pool in 8 hours. The empty pool needed to be filled as soon as possible, so the caretaker opened both pipes. However, he also mistakenly opened the drain. If the drain can empty the pool in 12 hours, how long will it take to fill the pool with both pipes and drain open?
----
large pipe rate: 1/6 job/hr
small pipe rate: 1/8 job/hr
drain rate: 1/12 job/hr
-------
together rate: 1/x job/hr
---
Equation:
rate + rate - rate = together rate
1/6 + 1/8 - 1/12 = 1/x
-----
Multiply thru by 24x to get:
4x + 3x - 2x = 24
5x = 24
x = 4.8 hrs (time to fill the pool)
=====================================
Cheers,
Stan H.